OpenStudy (anonymous):

The vertices of ∆ABC are A(-2, 2), B(6, 2), and C(0, 8). What is the perimeter of ∆ABC in units? MEDAL WILL BE GIVEN

jimthompson5910 (jim_thompson5910):

You would use the distance formula \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] to find the distance from A to B. This will give you the length of segment AB

jimthompson5910 (jim_thompson5910):

Do the same to find the lengths of BC and AC. Once you have those 3 side lengths, you add up the three lengths to get the perimeter.

OpenStudy (anonymous):

Okay thank you what about the area in square units?

OpenStudy (anonymous):

I didn't get it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

were you able to find the lengths of the sides?

OpenStudy (anonymous):